OpenStudy (anonymous):

Solve. 4cos²x + 4cosx + 1= 0

OpenStudy (lgbasallote):

let u = cosx 4u^2 + 4u + 1 = 0 (2u+1)(2u+1) 2u + 1 = 0 2u = -1 u = -1/2 cosx = -1/2 arc cosine it to get x

OpenStudy (callisto):

4cos²x + 4cosx + 1= 0 (2cosx+1)^2=0 (2cosx+1)=0 2cosx=-1 cosx = -1/2 x=120 degree or x=240 degree